During a certain year of the nominal interest rate was 8% the real interest rate was 3% and the CPI was 176.7 at the beginning of the year the cpi at the end of the year was
Answer #1

Real interest rate = Nominal interest rate- Inflation Rate

Inflation Rate = Nominal Interest Rate- Real interest rate

Inflation Rate = 8 - 3 = 5%

CPI End -Year - CPIPrevious-year Inflationr ate = C PIprevious-year

CPI - 176.7 176.7 End -Year

0.05*176.7 = CPI end of year - 176.7

CPI at the end of year = 8.835 + 176.7

CPI at the end of year = 185.535

Please contact if having any query thank you.